ONE Numbers is 6 more than another number. If you add 10 to 5 times the first number, the result is 6 times the second number.
Answer by Boreal(15235)

x is the smaller number, x+6 the larger number
5x+10=6(x+6)
5x+10=6x+36
x=-26
x+6=-20, the numbers
add 10 to 5 * -26 and it equals -120
That is 6 * (-20).